OpenStudy (anonymous):

The average life span of a phone battery is 21 months with a standard deviation of 1 months. Assuming the battery life is normally distributed, what is the probability that a battery chosen at random will last more than 24 months? Note: When finding Z-score round it to two decimal places and then use the Standard Normal Curve Table to find the area with four decimal places accuracy. Round the answer to four decimal places.

OpenStudy (anonymous):

Let X be the randomly chosen battery, then since the battery life is normally distributed: \[P(X>24)=P({X-21 \over 1}>{24-21 \over 1})=P(Z>3)=P(Z<-3)\]

OpenStudy (anonymous):

Now use the table to find P(Z<-3), you will get: P(Z<-3)=0.0013

OpenStudy (anonymous):

if your still on can you help me with this problem.... A company manufactures 57000 packages of jelly beans each week. On average, 2% of the packages do not seal properly. A random sample of 900 packages is selected at the end of the week. What is the probability that in this sample at least 22 are not properly sealed? Use a normal curve approximation for the binomial distribution.